OpenStudy (anonymous):

how many nonreal zeros does y=2x^3-x^2+10x-5 have

jhonyy9 (jhonyy9):

2x^3 -x^2 +10x -5=0 x^2(2x-1) +5(2x-1)=0 (x^2 +5)(2x -1)=0 x^2 +5=0 x^2 =-5 x_1,2=+/-sqrt(-5) =+/- isqrt5 so these will be the non real zeros 2x-1=0 2x=1 x=1/2 hope so much that is understandably good luck bye